OpenStudy (chrisneedshelp):

For the following system, use the second equation to make a substitution for y in the first equation. 2x + y = 5 y + 3 = 2x 2x + 2x - 3 = 5 2x + 3 - 2x = 5 2x + 2x + 3 = 5

OpenStudy (anonymous):

y=2x-3 2x+(2x-3)=5 4x=8 x=2

OpenStudy (chrisneedshelp):

so c?

OpenStudy (anonymous):

a. because 2(2)+2(2)-3=5 4+4-3=5 5=5

OpenStudy (chrisneedshelp):

thank you!